# Tapas are free! Training‑Free Adaptation of Programmatic Agents via LLM‑Guided Program Synthesis in Dynamic Environments

### v1

@@ Abstract
Autonomous agents in safety-critical applications must continuously adapt to dynamic conditions without compromising performance and reliability. This work introduces TAPA (Training-free Adaptation of Programmatic Agents), a novel framework that positions large language models (LLMs) as intelligent moderators of the symbolic action space. Unlike prior programmatic agents that typically generate a monolithic policy program or rely on fixed symbolic action sets, TAPA synthesizes and adapts modular programs for individual high-level actions, referred to as logical primitives. By decoupling strategic intent from execution, TAPA enables meta-agents to operate over an abstract, interpretable action space while the LLM dynamically generates, composes, and refines symbolic programs tailored to each primitive. Extensive experiments across cybersecurity and swarm intelligence domains validate TAPA’s effectiveness. In autonomous DDoS defense scenarios, TAPA achieves 77.7% network uptime while maintaining near-perfect detection accuracy in unknown dynamic environments. In swarm intelligence formation control under environmental and adversarial disturbances, TAPA consistently preserves consensus at runtime where baseline methods fail completely. This work promotes a paradigm shift for autonomous system design in evolving environments, from policy adaptation to dynamic action adaptation.
@@ 1 Introduction
Autonomous agents have become increasingly prevalent across critical domains such as cyber defense (Lohn et al. 2023), swarm intelligence control (Duan and Wang 2025), and autonomous driving (Kiran et al. 2021), where they are tasked with making timely and reliable decisions under complex conditions. Recent advancements in symbolic AI and reinforcement learning (RL) have greatly enhanced these agents’ ability to interact with their environments, enabling the acquisition of structured and abstract behaviors beyond pure rule-based control (Carnevali and Lippi 2024). The integration of neuro-symbolic approaches has emerged as the predominant paradigm for safety-critical applications, combining neural learning capabilities with symbolic rule reliability to overcome the black-box limitations of pure RL while maintaining the interpretability and formal guarantees required in critical domains (Marton et al. 2024). In such hybrid systems, symbolic modules are responsible for concrete execution by employing symbolic programs or rules as elements of the action space, enabling agents termed programmatic agents to operate through interpretable programmatic actions (Lyu et al. 2019; Verma et al. 2019). Compared to traditional actions, programmatic symbolic action spaces provide superior interpretability, verifiability, and integration of expert knowledge, making them well-suited for real-world applications (Lelis, Chen, and Sun 2024).
@@
However, these static agent frameworks fundamentally lack self-adaptive updating capabilities and struggle to match the continuous changed characteristic of dynamic operational environments. The rapid evolution of attack vectors, environmental conditions, and task requirements in safety-critical scenarios further exacerbates these limitations, resulting in degraded performance and rigid decision-making that overlooks the crucial need for adaptiveness to rapidly respond to unforeseen changes and urgent threats (Bhuyan et al. 2024; Li et al. 2024; Hu, Dong, and Huang 2025). Therefore, we summarized the existing approaches fall short in three aspects: (i) limited adaptability due to rigid symbolic action spaces; (ii) prohibitive retraining costs requiring extensive strategy overhaul or manual rule updates; and (iii) inadequate safeguards when outdated symbolic actions are applied to changed environments.
Although Large Language Models (LLMs) offer new possibilities with their powerful reasoning and generalization capabilities (Hadi et al. 2023; Guo et al. 2024b), current research largely focuses on end-to-end LLM decision agent, overlooking their potential to agents’ adaptiveness. Moreover, LLMs struggle in safety-critical, latency-sensitive environments, such as Distributed Denial of Service (DDoS) attack mitigation in cybersecurity scenarios where their inference overhead and the possibility of hallucinated output lead to inferior real-time performance compared to well-trained RL agents (Hager et al. 2024; Castro et al. 2025). This reveals a critical gap where existing autonomous systems lack adaptability, and LLMs lack efficiency and reliability.
Motivated by these complementary shortcomings, we adopt the paradigm shown in Figure 1 to illustrate our shift from coarse, policy‐level updates to fine‐grained action‐level synthesis, decomposing agent behavior into small, modular routines that can be synthesized and validated. We therefore ask: Can we utilize LLMs’ generative and reasoning capabilities for dynamic program synthesis and action-space moderation, enabling real-time adaptation without compromising the performance and reliability of expert-defined configurations or well-trained policies?
We answer this question with the TAPA framework (Figure 2). Our main contributions are:
  • Action-Space Moderation via LLMs: We introduce LLMs as dynamic moderators of an agent’s action space, on‐demand synthesizing, composing, and refining modular “logical‐primitive” programs to enable efficient, per‐action adaptation of autonomous systems.
  • TAPA (Training‐free Adaptation of Programmatic Agents): Based on this new moderator role, TAPA creates a RAG‐driven library of symbolic behaviors at design time and employs LLM‐guided shadow validation with hot‐swapping of synthesized programs at deployment to deliver real‐time responsiveness without any policy retraining.
  • Safety-Critical Empirical Validation: We evaluate TAPA in DDoS defense and multi-agent swarm formation under adversarial disturbances, demonstrating that its neuro-symbolic approach consistently outperforms both static symbolic and learning-based baselines.
@@ 2.1 Autonomous Decision Agents
Autonomous intelligent agents for real-time decision-making predominantly rely on traditional planning methods, reinforcement learning, or LLMs, each exhibiting critical limitations in dynamic environments requiring rapid adaptation. Traditional planning approaches such as STRIPS (Bylander 1994; Aineto, Jiménez, and Onaindia 2018), PDDL-based planners (Höller et al. 2020), and Hierarchical Task Networks (Hogg, Kuter, and Munoz-Avila 2009; Kelly, Botea, and Koenig 2008) require comprehensive domain specifications with predefined action sets, relying heavily on expert knowledge for domain modeling. While theoretically robust, their static nature and labor-intensive requirements significantly constrain practicality when rapid operational adaptations are required (Bhuyan et al. 2024).
Neuro-symbolic approaches integrate RL with symbolic AI to enhance capabilities through environmental interaction. These hybrid methods allow RL agents to operate over symbolic action spaces, or guide symbolic planning using learned value functions (Yang et al. 2018; Lyu et al. 2019; Shindo et al. 2025). However, they encounter critical limitations in adaptability under adversarial or highly dynamic conditions, where symbolic priors become obsolete and RL agents struggle to generalize across different operational scales due to rigid logic constraints (Hakim et al. 2025; Hu et al. 2025b). End-to-end LLM agents leverage natural language understanding to directly produce action sequences, enabling flexible reasoning and context-aware interaction (Yao et al. 2023). Despite their versatility, these agents face critical limitations in safety-critical deployments: inference latency hinders real-time responsiveness (Liang and Tong 2025), and outputs are prone to uncertainty and hallucination, undermining reliability even when guardrails are equipped (Dong et al. 2024a, b). In contrast, our TAPA framework positions LLMs as intelligent action space moderators, dynamically adapting action via programs synthesis. This approach addresses the adaptability-efficiency trade-off by enabling real-time symbolic program evolution without costly retraining, allowing agents to safely handle arbitrary scenarios including unexpected emergent situations through symbolic program adaptation.
@@ 2.2 Large Language Models for Code Generation
Recent advances in LLMs have revolutionized code generation capabilities, with models like GPT (Achiam et al. 2023), Claude (Anthropic 2024), and Code Llama (Roziere et al. 2023) demonstrating impressive performance on standard benchmarks such as HumanEval and MBPP (Jiang et al. 2024). These developments have led to widespread adoption of LLM-based coding assistants in software development workflows, including GitHub Copilot and various IDE integrations (El Haji, Brandt, and Zaidman 2024). However, code generation in safety-critical domains presents unique challenges that differentiate it from general-purpose programming tasks (Soroush et al. 2024). While RAG-based approaches (Parvez et al. 2021; Zhou et al. 2023) have enhanced code generation with external knowledge, the scarcity of high-quality, domain-specific code samples in safety-critical domains means RAG can only offer limited descriptive guidance, resulting in practically ineffective code generation. Recent work has explored various methods to address domain-specific code generation challenges, including specialized training datasets and domain-specific fine-tuning (Guo et al. 2024a; Zan et al. 2023; Hou et al. 2025). However, these approaches primarily focus on improving syntactic and functional correctness rather than ensuring domain-specific operational effectiveness in dynamic environments where code must adapt to rapidly evolving scenarios and threats. To enhance the practicality of generated programs, our framework constructs symbolic program pools through domain-specific multi-scenario simulation and employs provenance chains to guide LLM moderators in making informed decisions for symbolic program selection, composition, adaptation, and even new generation during runtime based on validated execution experiences.
@@ 3 Problem Formulation
where $P_{j}\in\mathcal{A}$ are selected symbolic programs, $\odot$ represents composition operations, and $\text{op}_{j}\in\{\land,\lor,+,-,\delta\}$ denotes logical operations including conjunction, disjunction, addition, deletion, and modification applied to program $P_{j}$. This formulation shifts adaptation from expensive policy retraining to efficient symbolic action space moderation, ensuring adaptive decision-making in dynamic scenarios.
@@ 4 TAPA Framework
All candidate programs must undergo validation through shadow simulation (Kuchta, Palikareva, and Cadar 2018; Microsoft 2023), a standard practice in industrial deployments where new algorithms run in parallel with production systems without affecting live operations. Within this shadow simulation environment, when significant environmental shifts are detected, the framework can also generate several alternative meta-policies $\{\pi_{\text{meta}}^{(1)},\pi_{\text{meta}}^{(2)},...,\pi_{\text{meta}}^{(k)}\}$ where $k$ represents the number of viable strategic alternatives, as backup options to maintain policy stability and ensure uninterrupted decision-making in safety and time-critical scenarios. This holistic design guarantees safety and performance before deployment while enabling quick strategic switches based on real-time feedback.
@@ 5 Experiments
To evaluate the effectiveness of TAPA across safety-critical domains, we conduct experiments on two representative cases: (1) Autonomous cyber defense, where agents must rapidly adapt to evolving DDoS attack patterns; and (2) Swarm formation control, where UAVs must maintain consensus under environmental disturbances and adversarial interference. These domains exemplify settings in which static symbolic systems struggle to adapt effectively (Dehghantanha, Yazdinejad, and Parizi 2023; Stodola, Nohel, and Horák 2025). Implementation details for both domains are provided in Appendix A. In addition, we perform ablation studies to validate the design rationale behind key components of the framework.
@@ 5.1 Autonomous Cyber Defense
This cybersecurity defense use case is designed to evaluate the effectiveness of action adaptation under evolving DDoS attack patterns and changing network conditions, and to assess the agent’s ability to maintain high detection accuracy, low false positives, and sustained network uptime (without retraining).
@@
We conduct our experiments using the NS3AI simulation platform (Yin et al. 2020), which combines high-fidelity network simulation (Henderson et al. 2008) with Python-based AI frameworks via shared memory linking, enabling real-time simulation and protocol-level traffic modeling. Our network topology mirrors standard enterprise architectures used in cyber defense testbeds (e.g., CIC-IDS2017, CIC-DDoS2019) (Sharafaldin et al. 2018), comprising a victim LAN with router/firewall, legitimate TCP clients, and a botnet launching high-rate attacks. To evaluate defense performance, we design five progressively complex scenarios (E2–E5) varying in attack sophistication, network size, and operational constraints, as summarized in Table 1. Following the NIST and MITRE cybersecurity configuration (Möller 2023), we model DDoS mitigation through four core defensive operations abstracted as logical primitives as shown in Example 1. We compare TAPA with several baseline methods, all of which are trained or configured in environment E1 and then evaluated in environments E2–E5 without prior exposure. These baselines operate with fixed action spaces consisting of one program per abstract intent defined in the example, denoted as $\mathcal{A}=\{P_{0,0},P_{1,0},P_{2,0},P_{3,0}\}$. In contrast, while TAPA also starts with the same initial program set $\mathcal{A}$, it activates action adaptation for the meta-agent when abnormal network performance is detected across consecutive time windows to optimize or replace the equipped programs.
Table 1: Environmental Evolution Scenarios Configuration
@@
  • Symbolic-Neural A Transformer-in-Transformer (TiT) based neural policy network (Mao et al. 2022).
@@
Overall, our TAPA framework introduces LLM generalization capability to facilitate action adaptation under evolving environments without over-reliance on perfectly predefined symbolic modules. The adaptive mechanism consistently maintains higher performance across dynamic cybersecurity environments through context-aware program synthesis and refinement, while the meta-policy enables rapid execution. Its balanced, neuro-symbolic architecture addresses the inherent limitations of static action spaces, constrained generalization capacity, and inefficient decision-making processes.
@@ 5.2 Formation Control of Swarm Intelligence
Formation control of swarm intelligence enables multiple autonomous agents (e.g., aircraft or robots) to maintain targeted formations through well-defined control algorithms. However, even carefully designed algorithms face significant limitations in dynamic environments with changing conditions, environmental disturbances, and adversarial threats.
@@
We evaluate TAPA in multi-aircraft formation control, where 10 aircraft maintain circular formation while adapting to environmental changes and adversarial threats. The baseline strategies follow Reynolds’ flocking rules (Reynolds 1987; Braga et al. 2017; Olfati-Saber and Murray 2004) and defensive strategies as implemented in (Hu et al. 2025a). Performance is evaluated across scenarios in Table 3 using formation radius and height metrics. We model formation control through two logical primitives: $L_{1}$ (Formation Control) for adjusting coordination parameters like cohesion distance and behavioral weights, and $L_{2}$ (Defend) for implementing defensive programs including outlier detection and noise injection for adversarial resilience. Each aircraft uses classical flocking behaviors (separation, cohesion, alignment, goal-seeking) through distributed decision-making within communication ranges, while TAPA enables real-time adjustment of formation variables including behavioral weights, distance thresholds, and defensive programs. When overall formation score based on radius and height continuously below threshold, TAPA’s LLM moderator analyzes environmental context and synthesizes appropriate parameter or program adaptations guided by expert knowledge and provenance chain experiences.
Table 3: Environmental Scenarios Configuration
@@
Figure 6 visualizes consensus achievement across scenarios, comparing baseline formation control, scenarios without TAPA, and TAPA-enabled adaptation through 3D trajectory visualization. Figure 5 provides quantitative analysis, illustrating TAPA’s consistent performance recovery across adversarial attacks and storm weather conditions. The results demonstrate that while baseline methods fail to maintain formation consensus under environmental changes, TAPA successfully triggers multiple action adaptations in dynamic environments, continuously adjusting programs and parameters to preserve aircraft consensus. TAPA’s ability to maintain near-baseline performance while adapting to diverse operational challenges validates symbolic program synthesis as an effective approach for safety-critical swarm intelligence applications requiring effectiveness, interpretability, and adaptiveness.

### v2

# Tapas Are Free! Training-Free Adaptation of Programmatic Agents via LLM-Guided Program Synthesis in Dynamic Environments

@@ Abstract
Autonomous agents in safety-critical applications must continuously adapt to dynamic conditions without compromising performance and reliability. This work introduces TAPA (Training-free Adaptation of Programmatic Agents), a novel framework that positions large language models (LLMs) as intelligent moderators of the symbolic action space. Unlike prior programmatic agents typically generate a monolithic policy program or rely on fixed symbolic action sets, TAPA synthesizes and adapts modular programs for individual high-level actions, referred to as logical primitives. By decoupling strategic intent from execution, TAPA enables meta-agents to operate over an abstract, interpretable action space while the LLM dynamically generates, composes, and refines symbolic programs tailored to each primitive. Extensive experiments across cybersecurity and swarm intelligence domains validate TAPA’s effectiveness. In autonomous DDoS defense scenarios, TAPA achieves 77.7% network uptime while maintaining near-perfect detection accuracy in unknown dynamic environments. In swarm intelligence formation control under environmental and adversarial disturbances, TAPA consistently preserves consensus at runtime where baseline methods fail. This work promotes a paradigm shift for autonomous system design in evolving environments, from policy adaptation to dynamic action adaptation.
@@ 1 Introduction
Autonomous agents have become increasingly prevalent across critical domains such as cyber defense (lohn2023autonomous), swarm intelligence control (duan2025enhancing), and autonomous driving (kiran2021deep), where they are tasked with making timely and reliable decisions under complex conditions. Recent advancements in symbolic AI and reinforcement learning (RL) have greatly enhanced these agents’ ability to interact with their environments, enabling the acquisition of structured and abstract behaviors beyond pure rule-based control (carnevali2024neuro). The integration of neuro-symbolic approaches has emerged as the predominant paradigm for safety-critical applications, combining neural learning capabilities with symbolic rule reliability to overcome the black-box limitations of pure RL while maintaining the interpretability and formal guarantees required in critical domains (marton2024sympol). In such hybrid systems, symbolic modules are responsible for concrete execution by employing symbolic programs or rules as elements of the action space, enabling agents termed programmatic agents to operate through interpretable programmatic actions (lyu2019sdrl; verma2019imitation). Compared to traditional actions, programmatic symbolic action spaces provide superior interpretability, verifiability, and integration of expert knowledge, making them well-suited for real-world applications (lelis2024programmatic).
@@
However, these static agent frameworks fundamentally lack self-adaptive updating capabilities and struggle to match the continuous changed characteristic of dynamic operational environments (hu2025stopreducingresponsibilityllmpowered). The rapid evolution of attack vectors, environmental conditions, and task requirements in safety-critical scenarios further exacerbates these limitations, resulting in degraded performance and rigid decision-making that overlooks the crucial need for adaptiveness to rapidly respond to unforeseen changes and urgent threats (bhuyan2024neuro; li2024logicity; hu2025trustorientedadaptiveguardrailslarge). Therefore, we summarized the existing approaches fall short in three aspects: (i) limited adaptability due to rigid symbolic action spaces; (ii) prohibitive retraining costs requiring extensive strategy overhaul or manual rule updates; and (iii) inadequate safeguards when outdated symbolic actions are applied to changed environments.
Although Large Language Models (LLMs) offer new possibilities with their powerful reasoning and generalization capabilities (hadi2023survey; guo2024large), current research largely focuses on end-to-end LLM decision agent, overlooking their potential to agents’ adaptiveness. Moreover, LLMs struggle in safety-critical, latency-sensitive environments, such as Distributed Denial of Service (DDoS) attack mitigation in cybersecurity scenarios where their inference overhead and the possibility of hallucinated output lead to inferior real-time performance compared to well-trained RL agents (hager2024evaluation; castro2025large). This reveals a critical gap where existing autonomous systems lack adaptability, and LLMs lack efficiency and reliability. Motivated by these complementary issues, we pursue a new paradigm shown in Figure 1: Can we harnesses LLMs’ generative and reasoning capabilities for dynamic program synthesis and action space moderation, enabling rapid adaptation while preserving efficiency and reliability of well-trained policies or expert-defined configurations?
To address this challenge, we propose the TAPA framework, illustrated in Figure 2. Overall, the key contributions of this paper are summarized as follows:
  • We introduce LLMs as intelligent action space moderators for autonomous agents, achieving a paradigm shift from costly policy to efficient action adaptation.
  • We propose TAPA, a training-free framework enabling agents to adapt to evolving environments through LLM-guided symbolic program synthesis while enhancing adaptability, interpretability and traceability.
  • We validate TAPA across safety-critical scenarios including cyber defense and swarm control, demonstrating superior performance in DDoS mitigation and formation control under adversarial disturbances.
@@ 2.1 Autonomous Decision Agents
Autonomous intelligent agents for real-time decision-making predominantly rely on traditional planning methods, reinforcement learning, or LLMs, each exhibiting critical limitations in dynamic environments requiring rapid adaptation. Traditional planning approaches such as STRIPS (bylander1994computational; aineto2018learning), PDDL-based planners (holler2020hddl), and Hierarchical Task Networks (hogg2009learning; kelly2008offline) require comprehensive domain specifications with predefined action sets, relying heavily on expert knowledge for domain modeling. While theoretically robust, their static nature and labor-intensive requirements significantly constrain practicality when rapid operational adaptations are required (bhuyan2024neuro).
Neuro-symbolic approaches integrate RL with symbolic AI to enhance capabilities through environmental interaction. These hybrid methods allow RL agents to operate over symbolic action spaces, or guide symbolic planning using learned value functions (yang2018peorl; lyu2019sdrl; shindo2025blendrl). However, they encounter critical limitations in adaptability under adversarial or highly dynamic conditions, where symbolic priors become obsolete and RL agents struggle to generalize across different operational scales due to rigid logic constraints (hakim2025ansr; 11077439). End-to-end LLM agents leverage natural language understanding to directly produce action sequences, enabling flexible reasoning and context-aware interaction (yao2023react). Despite their versatility, these agents face critical limitations in safety-critical deployments: inference latency hinders real-time responsiveness (liang2025llm), and outputs are prone to uncertainty and hallucination, undermining reliability even when guardrails are equipped (pmlr-v235-dong24c; dong2024safeguardinglargelanguagemodels). In contrast, our TAPA framework positions LLMs as intelligent action space moderators, dynamically adapting action via programs synthesis. This approach addresses the adaptability-efficiency trade-off by enabling real-time symbolic program evolution without costly retraining, allowing agents to safely handle arbitrary scenarios including unexpected emergent situations through symbolic program adaptation.
@@ 2.2 Large Language Models for Code Generation
Recent advances in LLMs have revolutionized code generation capabilities, with models like GPT (achiam2023gpt), Claude (claude3), and Code Llama (roziere2023code) demonstrating impressive performance on standard benchmarks such as HumanEval and MBPP (jiang2024survey). These developments have led to widespread adoption of LLM-based coding assistants in software development workflows, including GitHub Copilot and various IDE integrations (el2024using). However, code generation in safety-critical domains presents unique challenges that differentiate it from general-purpose programming tasks (soroush2024large). While RAG-based approaches (parvez-etal-2021-retrieval-augmented; zhou2023docprompting) have enhanced code generation with external knowledge, the scarcity of high-quality, domain-specific code samples in safety-critical domains means RAG can only offer limited descriptive guidance, resulting in practically ineffective code generation. Recent work has explored various methods to address domain-specific code generation challenges, including specialized training datasets and domain-specific fine-tuning (guo2024redcode; zan2023large; hou2025enhancing). However, these approaches primarily focus on improving syntactic and functional correctness rather than ensuring domain-specific operational effectiveness in dynamic environments where code must adapt to rapidly evolving scenarios and threats. To enhance the practicality of generated programs, our framework constructs symbolic program pools through domain-specific multi-scenario simulation and employs provenance chains to guide LLM moderators in making informed decisions for symbolic program selection, composition, adaptation, and even new generation during runtime based on validated execution experiences.
@@ 3 Problem Formulation
where $P_{j}\in\mathcal{A}$ are selected symbolic programs, $\odot$ represents composition operations that apply $\text{op}_{j}$ to programs, and $\text{op}_{j}\in\{\land,\lor,+,-,\delta\}$ denotes logical operations including conjunction, disjunction, addition, deletion, and modification applied to program $P_{j}$. This formulation shifts adaptation from expensive policy retraining to efficient symbolic action space moderation, ensuring adaptive decision-making in dynamic scenarios.
@@ 4 TAPA Framework
All candidate programs must undergo validation through shadow simulation (kuchta2018shadow; microsoft2023shadow), a standard practice in industrial deployments where new algorithms run in parallel with production systems without affecting live operations. Within this shadow simulation environment, when significant environmental shifts are detected, the framework can also generate several alternative meta-policies $\{\pi_{\text{meta}}^{(1)},\pi_{\text{meta}}^{(2)},...,\pi_{\text{meta}}^{(k)}\}$ where $k$ represents the number of viable strategic alternatives, as backup options to maintain policy stability and ensure uninterrupted decision-making in safety and time-critical scenarios. This holistic design guarantees safety and performance before deployment while enabling quick strategic switches based on real-time feedback.
@@ 5 Experiments
To evaluate the effectiveness of TAPA across safety-critical domains, we conduct experiments on two representative cases: (1) Autonomous cyber defense, where agents must rapidly adapt to evolving DDoS attack patterns; and (2) Swarm formation control, where UAVs must maintain consensus under environmental disturbances and adversarial interference. These domains exemplify settings in which static symbolic systems struggle to adapt effectively (dehghantanha2023autonomous; stodola2025dynamic). Our results demonstrate that TAPA enables training-free action adaptation while preserving interpretability and operational efficiency. Implementation details for both domains are provided in Appendix A. In addition, we perform ablation studies to validate the design rationale behind key components of the framework.
@@ 5.1 Autonomous Cyber Defense
This cybersecurity defense use case is designed to evaluate the effectiveness of action adaptation under evolving DDoS attack patterns and changing network conditions, and to assess the agent’s ability to maintain high detection accuracy, low false positives, and sustained network uptime (without policy retraining, which typically needs hours to adapt to unknown environments).
@@
We conduct our experiments using the NS3AI simulation platform (yin2020ns3), which combines high-fidelity network simulation (henderson2008network) with Python-based AI frameworks via shared memory linking, enabling real-time simulation and protocol-level traffic modeling. Our network topology mirrors standard enterprise architectures used in cyber defense testbeds (e.g., CIC-IDS2017, CIC-DDoS2019) (sharafaldin2018toward), comprising a victim LAN with router/firewall, legitimate TCP clients, and a botnet launching high-rate attacks. To evaluate defense performance, we design five progressively complex scenarios (E2–E5) varying in attack sophistication, network size, and operational constraints, as summarized in Table 1. Following the NIST and MITRE cybersecurity configuration (moller2023nist), we model DDoS mitigation through four core defensive operations abstracted as logical primitives as shown in Example 1. We compare TAPA with several baseline methods, all of which are trained or configured in environment E1 and then evaluated in environments E2–E5 without prior exposure. These baselines operate with fixed action spaces consisting of one program per abstract intent defined in the example, denoted as $\mathcal{A}=\{P_{0,0},P_{1,0},P_{2,0},P_{3,0}\}$. In contrast, while TAPA also starts with the same initial program set $\mathcal{A}$, it activates action adaptation for the meta-agent when abnormal network performance is detected across consecutive time windows to optimize or replace the equipped programs.
@@
  • Symbolic-Neural A Transformer-in-Transformer (TiT) based neural policy network (mao2022transformer).
@@
Overall, our TAPA framework introduces LLM generalization capability to facilitate action adaptation under evolving environments without over-reliance on perfectly predefined symbolic modules. The adaptive mechanism consistently maintains higher performance across dynamic cybersecurity environments through context-aware program synthesis and refinement, while the meta-policy enables rapid execution. Its balanced, neuro-symbolic paradigm addresses the inherent limitations of static action spaces, constrained generalization capacity, and inefficient decision-making processes, providing a principled and scalable solution for training-free autonomous adaptation in safety-critical dynamic environments.
@@ 5.2 Formation Control of Swarm Intelligence
Formation control of swarm intelligence enables multiple autonomous agents (e.g., aircraft or robots) to maintain targeted formations through well-defined control algorithms. However, even carefully designed algorithms face significant limitations in dynamic environments with changing conditions, environmental disturbances, and adversarial threats. TAPA addresses these challenges by enabling real-time program synthesis for swarm coordination, dynamically adapting control parameters and logic. We evaluate TAPA against baseline methods across diverse scenarios, demonstrating its effectiveness in maintaining formation coherence under varying operational conditions.
@@
We evaluate TAPA in multi-aircraft formation control, where 10 aircraft maintain circular formation while adapting to environmental changes and adversarial threats. The baseline strategies follow Reynolds’ flocking rules (reynolds1987flocks; braga2017collision; olfati2004consensus) and defensive strategies as implemented in (hu2025enhancingrobustnessllmdrivenmultiagent). Performance is evaluated across scenarios in Table 3 using formation radius and height metrics. We model formation control through two logical primitives: $L_{1}$ (Formation Control) for adjusting coordination parameters like cohesion distance and behavioral weights, and $L_{2}$ (Defend) for implementing defensive programs including outlier detection and noise injection for adversarial resilience. Each aircraft uses classical flocking behaviors (separation, cohesion, alignment, goal-seeking) through distributed decision-making within communication ranges, while TAPA enables real-time adjustment of formation variables including behavioral weights, distance thresholds, and defensive programs. When overall formation score based on radius and height continuously below threshold, TAPA’s LLM moderator analyzes environmental context and synthesizes appropriate parameter or program adaptations guided by expert knowledge and provenance chain experiences.
@@
Figure 6 and Appendix C visualizes consensus achievement across scenarios, comparing baseline formation control, scenarios without TAPA, and TAPA-enabled adaptation through 3D trajectory visualization. Figure 5 provides quantitative analysis, illustrating TAPA’s consistent performance recovery across adversarial attacks and storm weather conditions. The results demonstrate that while baselines fail to maintain formation consensus under environmental changes, TAPA successfully triggers multiple action adaptations in dynamic environments, continuously adjusting programs and parameters to preserve aircraft consensus. TAPA’s ability to maintain near-baseline performance while adapting to diverse operational challenges validates symbolic program synthesis as an effective approach for safety-critical swarm intelligence applications requiring effectiveness, interpretability, and adaptiveness.